OpenStudy (anonymous):

The map of a walking trail is drawn on a coordinate grid with three points of interest. The trail starts at R(−2, 4) and goes to S(3, 4) and continues to T(3, −1). The total length of the walking trail is ____ units. (Input whole numbers only.)

OpenStudy (anonymous):

new qustion posted

OpenStudy (kittiwitti1):

gotcha. okay:

OpenStudy (kittiwitti1):

 RS=x, ST=y

OpenStudy (kittiwitti1):

RS only changes x coordinates so x= 3-(-2)=? ST only changes y coordinates so y= 4-(-1)=? add x and y: x+y = ?

OpenStudy (kittiwitti1):

x+y = [3-(-2)]+[4-(-1)]=?

OpenStudy (anonymous):

thank you

OpenStudy (kittiwitti1):

np